OpenStudy (anonymous):

The geometric sequence that tells the perimeter of each stage in the building of Koch's snowflake has a first term of 3 and a ratio of 4/3. At the 18th stage in the construction of this snowflake, the length of each tiny side would be about the size of an atom! Using the explicit formula for the nth term of a geometric sequence, what is the length of the perimeter a18 to the nearest whole number? A. 399 B. 299 C. 200 D. 133 E. 99

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

@SolomonZelman can you help?

OpenStudy (solomonzelman):

\(\large\color{slate}{ a_1=3 }\) \(\large\color{slate}{ r=4/3 }\) \(\large\color{slate}{ a_{n}=a_1\times(r)^{n-1} }\) \(\large\color{slate}{ \Downarrow }\) \(\large\color{slate}{ a_{18}=3\times(4/3)^{18-1} }\)

OpenStudy (solomonzelman):

I got to go soon, sorry

OpenStudy (anonymous):

@dtan5457 can you help me?

OpenStudy (anonymous):

@magepker728

OpenStudy (anonymous):

Someone help.. I am still confused on this

OpenStudy (anonymous):

where are you confused solomonzelman has put you everthing together, and all u have to do is to solve it

OpenStudy (anonymous):

I put it in a calcultor to make it go quicker and it like gives me some weird numbers that don't come close to the answers.. and im lost. P.s. really slow with math so..

OpenStudy (anonymous):

ok let take it step by

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):